A long solenoid has a length of 0.65m and contains 1400 turns of a wire. There is a current of 4.7A in the wire. What is the magnitude of the magnetic field within the solenoid?
Answer #1

The Magnitude of uniform magnetic field inside the solenoid, B = where = 4π(10^-7)

n= number of turns per unit length= N/L

I = current flowing

Given n=1400/0.65= 2153.84 m^-1

I = 4.7 A

Therefore B= 4π (10^-7)* (2153.84)* 4.7= 1.27 x (10^-2) Tesla